# Disentangling Hippocampal Shape Variations: A Study of Neurological Disorders Using Graph Variational Autoencoder with Contrastive Learning

### v1

@@ Abstract
This paper presents a comprehensive study focused on disentangling hippocampal shape variations from diffusion tensor imaging (DTI) datasets within the context of neurological disorders. Leveraging a Graph Variational Autoencoder (VAE) enhanced with Supervised Contrastive Learning, our approach aims to improve interpretability by disentangling two distinct latent variables corresponding to age and the presence of diseases. In our ablation study, we investigate a range of VAE architectures and contrastive loss functions, showcasing the enhanced disentanglement capabilities of our approach. This evaluation uses synthetic 3D torus mesh data and real 3D hippocampal mesh datasets derived from the DTI hippocampal dataset. Our supervised disentanglement model outperforms several state-of-the-art (SOTA) methods like attribute and guided VAEs in terms of disentanglement scores. Our model distinguishes between age groups and disease status in patients with Multiple Sclerosis (MS) using the hippocampus data. Our Graph VAE with Supervised Contrastive Learning shows the volume changes of the hippocampus of MS populations at different ages, and the result is consistent with the current neuroimaging literature. This research provides valuable insights into the relationship between neurological disorder and hippocampal shape changes in different age groups of MS populations using a Graph VAE with Supervised Contrastive loss.
@@
Keywords: Disentangled Representation Learning, Graph Variational Autoencoder, Deep Learning, Contrastive Learning, Neurological Disorders, Medical Imaging, Hippocampal Shape Variations, Diffusion Tensor Image.
@@ 1.1 Motivation and Objective
We employ graph VAE to obtain interpretable latent dimensions and generate valid shapes similar to the training dataset. We use a modified contrastive loss (Frosst et al., 2019) as a latent space disentanglement strategy to isolate two data generative factors: age and disease (MS). Age is represented by continuous values, while disease is labeled by discrete values. This combination of regression and classification tasks is frequently encountered in medical imaging. Our model formulation results in more interpretable latent codes and enables control over the generative process based on the specified factors (both classification and regression). As part of our validation process for the proposed method, we develop a 3D synthetic torus dataset with four factors of variability. During training and testing, we disentangled two of these factors using labels. Our results demonstrate supervised disentanglement using both classification and regression data, both combined and separately.
@@ 1.2 Related Works
The disentanglement of the latent space has been the focus of numerous research works. In this section, we provide an overview of both supervised and unsupervised methods for disentangling latent variables. In section 1.2.1, we discuss the vanilla VAE and related works that enhance the disentanglement performance of the vanilla VAE. Most of these methods are unsupervised and disentangle the latent space without any prior knowledge of which variable disentangles which data-generating factor. We also examine supervised VAEs that disentangle specific latent variables using the data labels, and these methods exhibit strong disentangling performance for specific factors when compared to unsupervised approaches. Furthermore, we present some contrastive learning-based methods that enhance disentanglement, although they are also unsupervised methods. Finally, in section 1.2.2, we explore disentanglement techniques that employ 3D mesh data with supervision and are related to our proposed method.
@@ 1.2.1 Disentangled Latent Representation VAEs
1.2.2 Disentanglement in 3D Mesh Data using Graph VAEs
@@
While most disentangling methods are designed for images, researchers also use supervised VAEs to disentangle specific attributes in 3D mesh datasets. One of the research (Foti et al., 2022) introduced a self-supervised method for training a 3D shape VAE aimed at achieving a disentangled latent representation of identity features in 3D generative models for faces and bodies. The approach involves mini-batch feature swapping between various shapes to optimize mini-batch generation and formulate a loss function based on known differences and similarities in latent representations. Sun et al. (2022) proposed a VAE framework to disentangle identity and expression from 3D input faces that have a wide variety of expressions.
@@
The two approaches mentioned for 3D VAE are not applicable to our specific medical domain problem, as we find neither feature swapping nor unsupervised learning appropriate. In our case, we possess labeled data and aim to disentangle multiple latent variables (for classification and regression) with supervision because supervised training increases the disentangling and reconstruction performance according to the previous research we discussed. Our work partially uses the process proposed by Kiechle et al. (2023), who explore a supervised variational graph autoencoder to understand and explain the variability in anatomical shapes. However, they only use the excitation-inhibition mechanism for the regression problem and used two additional neural networks for their method, which is different from our proposed method.
@@
The domain of medical imaging often necessitates the disentanglement of various factors, including but not limited to age, diseases, and gender, through both classification and regression techniques. Therefore, we focus on simultaneous classification and regression techniques in the medical imaging domain with better loss functions. To the best of our knowledge, no prior work has utilized contrastive loss for both the classification and regression tasks (simultaneously) with a guided graph VAEs to disentangle multiple latent variables. Our proposed method demonstrates superior disentanglement compared to guided VAE (Ding et al., 2020) and attribute VAE (Cetin et al., 2023) while achieving comparable generative quality and speed.
@@ 2 Materials and Methods
Our proposed VAE designed for deep graph convolution operates with an input comprising 3D mesh vertices denoted as $X=[x_{0},x_{1},...,x_{N-1}]^{T}\in\mathbb{R}^{N\times F}$, where $F$ represents the feature dimension, and $N$ is the total number of vertices per mesh. In the context of 3D mesh data, $F$ is specified as 3, and $X$ contains the coordinates of each vertex.
@@
Figure 1: Overall architecture of our method. We have graph VAE with an encoder $f_{\phi}(x)$ and decoder $f_{\theta}(z)$ where $x$ is the input 3D mesh and $z$ represents the latent space. $L_{vae}$ represents the VAE loss that combines reconstruction and KL divergence loss. Another two losses are classification loss $L_{contr}^{cls}$ and regression loss $L_{contr}^{reg}$, where a specific latent variable is disentangled for a specific feature (continuous or discrete). We use the first variable for contrastive classification loss ($z_{1}$ corresponds to binary labels, and the rest variables are uncorrelated to the labels). The second variable $z_{2}$ corresponds to regression loss, and the rest variables are uncorrelated to the continuous labels.
@@ 2.3 Supervised Contrastive VAE (Ours)
$L_{contr}^{reg}=-\frac{1}{b}\sum_{i\in 1..b}\log\left(\frac{\sum_{\begin{subarray}{c}j\in 1..b\\ j\neq i\\ |y_{i}-y_{j}|\leq Th\end{subarray}}e^{-\frac{||z_{2}^{i}-z_{2}^{j}||^{2}}{T}}}{\lambda_{1}\sum_{\begin{subarray}{c}k\in 1..b\\ k\neq i\end{subarray}}e^{-\frac{||z_{2}^{i}-z_{2}^{k}||^{2}}{T}}+\lambda_{2}\sum_{\begin{subarray}{c}j\in 1..b\\ j\neq i\\ |y_{i}-y_{j}|\leq Th\end{subarray}}e^{-\frac{\sum_{1,3..d_{z}}||z_{d}^{i}-z_{d}^{j}||^{2}}{(d-1)T}}}\right)$ (10)
@@ 3.4.1 Comparison
We compare the models in terms of disentanglement score (SAP), the correlation between the latent variables and labels (Corr.), accuracy (Acc.), and MSE score in predicting the labels from the values of the latent variables using the K-nearest neighbor classifier. The reported SAP scores in table 1 are calculated by averaging the SAP scores for $z_{0}$ and $z_{1}$ variables and the models are trained simultaneously for classification and regression tasks for all the scores.
@@ 3.4.5 MS vs Hippocampal Volume Across Age Groups
Figure 6: Volume differences between healthy and MS hippocampus are shown using six age ranges. Volume difference for the left and right hippocampus is shown separately. The graph depicts a higher right hippocampus volume difference across ages. The volumes are calculated by generating 10 shapes (using our model) from each range, and then taking the average.
@@ 4.2 Conclusion
In this paper, we propose a novel approach for disentangling 3D mesh shape (hippocampal or synthetic) variations from DTI or synthetic datasets and applied in the context of neurological disorders. Our method, which uses a Graph VAE enhanced with Supervised Contrastive Learning, exhibits superior disentanglement capabilities, particularly in identifying age and disease status in patients with MS. Additionally, our method demonstrates comparable or better performance in all other metrics. The validity of our method is also demonstrated by a synthetic torus dataset.

### v2

# Disentangling Hippocampal Shape Variations: A Study of Neurological Disorders Using Mesh Variational Autoencoder with Contrastive Learning

@@ Abstract
This paper presents a comprehensive study focused on disentangling hippocampal shape variations from diffusion tensor imaging (DTI) datasets within the context of neurological disorders. Leveraging a Mesh Variational Autoencoder (VAE) enhanced with Supervised Contrastive Learning, our approach aims to improve interpretability by disentangling two distinct latent variables corresponding to age and the presence of diseases. In our ablation study, we investigate a range of VAE architectures and contrastive loss functions, showcasing the enhanced disentanglement capabilities of our approach. This evaluation uses synthetic 3D torus mesh data and real 3D hippocampal mesh datasets derived from the DTI hippocampal dataset. Our supervised disentanglement model outperforms several state-of-the-art (SOTA) methods like attribute and guided VAEs in terms of disentanglement scores. Our model distinguishes between age groups and disease status in patients with Multiple Sclerosis (MS) using the hippocampus data. Our Mesh VAE with Supervised Contrastive Learning shows the volume changes of the hippocampus of MS populations at different ages, and the result is consistent with the current neuroimaging literature. This research provides valuable insights into the relationship between neurological disorder and hippocampal shape changes in different age groups of MS populations using a Mesh VAE with Supervised Contrastive loss.
@@
Keywords: Disentangled Representation Learning, Mesh Variational Autoencoder, Deep Learning, Contrastive Learning, Neurological Disorders, Medical Imaging, Hippocampal Shape Variations, Diffusion Tensor Image.
@@ 1.1 Motivation and Objective
We employ Mesh VAE to obtain interpretable latent dimensions and generate valid shapes similar to the training dataset. We use a modified contrastive loss (Frosst et al., 2019) as a latent space disentanglement strategy to isolate two data generative factors: age and disease (MS). Age is represented by continuous values, while disease is labeled by discrete values. This combination of regression and classification tasks is frequently encountered in medical imaging. Our model formulation results in more interpretable latent codes and enables control over the generative process based on the specified factors (both classification and regression). As part of our validation process for the proposed method, we develop a 3D synthetic torus dataset with four factors of variability. During training and testing, we disentangled two of these factors using labels. Our results demonstrate supervised disentanglement using both classification and regression data, both combined and separately.
@@ 1.2 Related Works
The disentanglement of the latent space has been the focus of numerous research works. In this section, we provide an overview of both supervised and unsupervised methods for disentangling latent variables. In section 1.2.1, we discuss the vanilla VAE and related works that enhance the disentanglement performance of the vanilla VAE. Most of these methods are unsupervised and disentangle the latent space without any prior knowledge of which variable disentangles which data-generating factor. We also examine supervised VAEs that disentangle specific latent variables using the data labels, and these methods exhibit strong disentangling performance for specific factors when compared to unsupervised approaches. Furthermore, we present some contrastive learning-based methods that enhance disentanglement, although they are also unsupervised methods. Additionally, we explore some graph autoencoder methods. Finally, in section 1.2.2, we explore disentanglement techniques that employ 3D mesh data with self-supervision and conditional VAEs and are related to our proposed method.
@@ 1.2.1 Disentangled Latent Representation VAEs
Another type of autoencoder called Graph autoencoder has become a crucial tool for studying graph-structured data. It enables the learning of meaningful representations for tasks such as node classification, link prediction, and clustering. A pioneering contribution to this domain was made by Kipf et al., who introduced the Graph Convolutional Network (GCN) in an autoencoder framework. Their approach, known as the Variational Graph Autoencoder (VGAE) (Kipf and Welling, 2016), combines the power of GCNs to aggregate and propagate node features with a variational autoencoder’s capacity for learning interpretable latent representations. Further advancements include the work of Pan et al.’s Adversarially Regularized Graph Autoencoder (ARGA), and Adversarially Regularized Variational Graph autoencoder (ARVGA) (Pan et al., 2018), which incorporates adversarial training to enhance the robustness of learned representations, and Wang et al.’s Marginalized Graph Autoencoder (MGAE) (Wang et al., 2017), which introduced a denoising-based approach specifically tailored for graph clustering tasks.
1.2.2 Disentanglement in 3D Mesh Data using Mesh VAEs
@@
While most disentangling methods are designed for images, researchers also use self-supervised and conditional VAEs to disentangle specific attributes in 3D mesh datasets. One of the research (Foti et al., 2022) introduced a self-supervised method for training a 3D shape VAE aimed at achieving a disentangled latent representation of identity features in 3D generative models for faces and bodies. The approach involves mini-batch feature swapping between various shapes to optimize mini-batch generation and formulate a loss function based on known differences and similarities in latent representations. Sun et al. (2022) proposed a VAE framework to disentangle identity and expression from 3D input faces that have a wide variety of expressions.
@@
The two approaches mentioned for 3D VAE are not applicable to our specific medical domain problem, as we find neither feature swapping nor unsupervised learning appropriate. In our case, we possess labeled data and aim to disentangle multiple latent variables (for classification and regression) with supervision because supervised training increases the disentangling and reconstruction performance according to the previous research we discussed. Our work partially uses the process proposed by Kiechle et al. (2023), who explore a supervised variational Mesh autoencoder to understand and explain the variability in anatomical shapes. However, they only use the excitation-inhibition mechanism for the regression problem and used two additional neural networks for their method, which is different from our proposed method.
@@
The domain of medical imaging often necessitates the disentanglement of various factors, including but not limited to age, diseases, and gender, through both classification and regression techniques. Therefore, we focus on simultaneous classification and regression techniques in the medical imaging domain with better loss functions. Our decision to focus on the two latent factors (classification and regression tasks) in the hippocampal study was primarily guided by their strong biological relevance and interpretability, particularly in the context of neurodegenerative diseases. These factors align with well-understood morphological variations in hippocampal structures, making our model’s output more relevant for clinical research. The availability of supervised labels for these specific dimensions further supports our choice, allowing us to validate the model’s disentanglement and ensure that the learned representations are meaningful and clinically significant. To the best of our knowledge, no prior work has utilized contrastive loss for both the classification and regression tasks (simultaneously) with a guided mesh VAEs to disentangle multiple latent variables. Our proposed method demonstrates superior disentanglement compared to guided VAE (Ding et al., 2020) and attribute VAE (Cetin et al., 2023) while achieving comparable generative quality and speed.
@@ 2 Materials and Methods
Our proposed VAE designed for deep mesh convolution operates with an input comprising 3D mesh vertices denoted as $X=[x_{0},x_{1},...,x_{N-1}]^{T}\in\mathbb{R}^{N\times F}$, where $F$ represents the feature dimension, and $N$ is the total number of vertices per mesh. In the context of 3D mesh data, $F$ is specified as 3, and $X$ contains the coordinates of each vertex.
@@
Our method uses SpiralNet++ to exploit the local geometric structure of mesh data, preserving spatial relationships between vertices, which is crucial for accurately modeling the hippocampus and torus data. In contrast, we do not use PointNet-type models that focus on global features and may lose critical local geometric information. SpiralNet++ ensures consistent capture of local features through a fixed template, aligning with our goal of disentangling specific factors within hippocampus data (Gong et al., 2019). The importance of maintaining local mesh structure for accurate shape analysis, as highlighted by Litany et al. (Litany et al., 2018), further justifies the use of SpiralNet++ over PointNet-type models.
@@
Figure 1: Overall architecture of our method. We have mesh VAE with an encoder $f_{\phi}(x)$ and decoder $f_{\theta}(z)$ where $x$ is the input 3D mesh and $z$ represents the latent space. $L_{vae}$ represents the VAE loss that combines reconstruction and KL divergence loss. Another two losses are classification loss $L_{contr}^{cls}$ and regression loss $L_{contr}^{reg}$, where a specific latent variable is disentangled for a specific feature (continuous or discrete). We use the first variable for contrastive classification loss ($z_{1}$ corresponds to binary labels, and the rest variables are uncorrelated to the labels). The second variable $z_{2}$ corresponds to regression loss, and the rest variables are uncorrelated to the continuous labels.
@@ 2.3 Supervised Contrastive VAE (Ours)
$L_{contr}^{reg}=-\frac{1}{b}\sum_{i\in 1..b}\log\left(\frac{\sum_{\begin{subarray}{c}j\in 1..b\\ j\neq i\\ |y_{i}-y_{j}|\leq Th\end{subarray}}e^{-\frac{||z_{2}^{i}-z_{2}^{j}||^{2}}{T}}}{\lambda_{1}\sum_{\begin{subarray}{c}k\in 1..b\\ k\neq i\end{subarray}}e^{-\frac{||z_{2}^{i}-z_{2}^{k}||^{2}}{T}}+\lambda_{2}\sum_{\begin{subarray}{c}j\in 1..b\\ j\neq i\\ |y_{i}-y_{j}|\leq Th\end{subarray}}e^{-\frac{\sum_{d\in{1,3..d_{z}}}||z_{d}^{i}-z_{d}^{j}||^{2}}{(d-1)T}}}\right)$ (10)
@@ 3.4.1 Comparison
We compare the models in terms of disentanglement score (SAP), the correlation between the latent variables and labels (Corr.), accuracy (Acc.), and MSE score in predicting the labels from the values of the latent variables using the K-nearest neighbor classifier. The reported SAP scores in table 1 are calculated by averaging the SAP scores for $z_{1}$ and $z_{2}$ variables and the models are trained simultaneously for classification and regression tasks for all the scores.
@@ 3.4.2 Ablation Study
Figure 4: Effect of Spiral Sequence Length on SAP Score for our Torus and Hippocampus Datasets.
Additionally, we conducted an ablation study to justify the selection of a mesh-based convolutional neural network.The plot in figure 4 demonstrates that varying the spiral sequence length affects the SAP score for both the hippocampus and torus datasets. The non-linear relationship between spiral length and SAP scores shows that optimal performance occurs at specific lengths, implying that neighborhood connectivity impacts disentanglement. This justifies our use of SpiralNet++, which leverages such connectivity effectively.
@@ 3.4.5 MS vs Hippocampal Volume Across Age Groups
Figure 7: Volume differences between healthy and MS hippocampus are shown using six age ranges. Volume differences for the left and right hippocampus are shown separately for all three supervised disentanglement methods (SG VAE, Attribute VAE, and our SC VAE). The plot depicts a higher right hippocampus volume difference across ages derived by all three methods. However, our method shows the highest volume difference, which is closer to the average trend found in (Valdés Cabrera et al., 2023). The volumes are calculated by generating 10 shapes (using the three models) from each range, and then taking the average.
@@ 4.2 Conclusion
In this paper, we propose a novel approach for disentangling 3D mesh shape (hippocampal or synthetic) variations from DTI or synthetic datasets and applied in the context of neurological disorders. Our method, which uses a Mesh VAE enhanced with Supervised Contrastive Learning, exhibits superior disentanglement capabilities, particularly in identifying age and disease status in patients with MS. Additionally, our method demonstrates comparable or better performance in all other metrics. The validity of our method is also demonstrated by a synthetic torus dataset.